A 0.100g sample of a compound C,H, and O is burned in oxygen producing 0.1783g of CO2 and 0.0734g of H2O. Determine the empirical formula of the compound, listing elements in the order of CnHnOn


1
Expert's answer
2020-10-09T13:45:30-0400

Solution:

We need to find out the percent composition of hydrogen in water (H2O) and of carbon in carbon dioxide (CO2). This will allow to find out how many grams of hydrogen and carbon, respectively, were produced by the combustion reaction.


One mole of water (H2O) contains 2 moles of hydrogen.

The percent composition of hydrogen in water is:

(2 × 1.00794 g/mol) / (18.015 g/mol) = 0.1119 or 11.19%

This means that every 100 grams of water contain 11.19 grams of hydrogen.

In our case, a 0.0734 g sample of water contains:

0.0734 g H2O × (11.19 g H / 100 g H2O) = 0.00821 g H


One mole of carbon dioxide (CO2) contains one mole of carbon.

The percent composition of carbon in carbon dioxide is:

(1 × 12.011 g/mol) / (44.01 g/mol) = 0.2729 or 27.29%

This means that every 100 grams of carbon dioxide contain 27.29 grams of carbon.

In our case, a 0.1783 g sample of carbon dioxide contains:

0.1783 g CO2 × (27.29 g C / 100 g CO2) = 0.04866 g C


The initial mass of the sample is equal to 0.100 g.

This suggests that the sample also contains:

moxygen = msample - mhydrogen - mcarbon

moxygen = 0.100 g - 0.00821 g - 0.04866 g = 0.04313 g


Using the molar masses and masses of each elements to convert their to moles:

For C: 0.04866 g / 12.011 g mol-1 = 0.00405 moles C

For H: 0.00821 g / 1.00794 g mol-1 = 0.008145 moles H

For O: 0.04313 g / 15.9994 g mol-1 = 0.002696 moles O


Coefficients for the tentative empirical formula are derived by dividing each molar amount by the lesser of the three:

0.00405 moles C / 0.002696 = 1.5

0.008145 moles H / 0.002696 = 3.0

0.002696 moles O / 0.002696 = 1.0


This means that: C1.5H3O1


Since we are looking for the smallest whole number ratio that exists between the elements, multiply each subscript by 2 to get:

(C1.5H3O1)2 ⇒ C3H6O2


Answer: The empirical formula of the compound is C3H6O2